40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5120     Entry time: Fri Aug 5 02:00:09 2011
Author: kiwamu 
Type: Summary 
Category: General 
Subject: summary of today's invac work 

[Jamie, Jenne, Suresh, Steve, Koji, Kiwamu]

We got two green beams coming out from the chambers !

 

Summary of today's invac work :

    - removed the access connector and the BS north door

    - realigned the X and Y arm to the green beams.

    - installed a HWP on the ETMY table to rotate the polarization of the green beam to P.

    - repositioned the first periscope on the BS table.

    - repositioned the second periscope on the IOO table.

    - steered some green mirrors on the IOO and OMC chamber to let the Y green beam come out to the PSL table.

    - installed a PBS in front of the first periscope to spatially overwrap two green beams.

    - adjusted the incident angle of the PBS to maximize the power of the Y green beam, which is transmitted through it.

   - steer two mirrors on the BS table to align the X green beam

   - installed two beam dumps, one is near the PBS to eliminate a ghost in the X green beam, and the other is on the back side of IPPOS/ANG pick off window.

   - closed the doors.

 

Some Notes:

 When steering the final green mirror on the OMC table, accidentally we changed the alignment of the MC incident mirror.

So the alignment of the incident beam going into MC has changed, and we haven't re-aligned it yet.

 During we were installing the PBS on the BS table, we found that the allowable incident angle for the Y beam is about ~ 55 deg, which maximizes the amount of the transmitted Y green.

Since the PBS had been considered to be 45 deg incident in our optical layout, this required several modifications in the green mirrors.

To have a clear X green beam path going into the PBS, we had to slide the PBS and periscope to the West.

The periscope is now sitting on the very edge of the BS table, and in fact ~ 20% of the bottom plate of the periscope is already sticking out.

Also since 30% of the area of the PBS's post is on a hole, which is somehow for the stack, we had to use three dog clamps instead of a folk clamps to make the contact tight.

ELOG V3.1.3-